OpenStudy (anonymous):

The chart below shows the number of teeth of several snakes Anaconda Python Coral Garter Boa 100 112 76 36 How many teeth does the Boa snake have if the average number of teeth for the 5 snakes is 88?

OpenStudy (anonymous):

Do you need the work or just the answer?

OpenStudy (anonymous):

The Boa should have 116 teeth

OpenStudy (anonymous):

I got it know thank you I got 116:)

OpenStudy (anonymous):

No problem!

OpenStudy (dwelsher):

(100+112+76+36+x)/5=88 (332+x/)5=88 x=108